# Universal non-Hermitian transport in disordered systems

### v1

@@ Abstract
where $g=v\eta^{2}/D^{4}$, $K_{d}=S_{d}/(2\pi)^{d}$ with $S_{d}$ being the solid angle in $d$ dimension, and the rescaling $\mathbf{x}\rightarrow e^{l}\mathbf{x}$ is performed. The flow equation yields two fixed points, $g_{1}^{\ast}=0$, $g^{\ast}_{2}=d(d-4)/[2(12-5d)K_{d}]$. Notably, $g_{1}^{\ast}$ is unstable for $d<4$, while $g^{\ast}_{2}$ is physically untenable for $d<12/5$ due to its negative value, conflicting with the definition. Thus, in 1D and 2D, the system does not transition to new phases upon rescaling, validating our analysis for weak disorders. However, for $d=3$, $g^{\ast}_{2}=1/(2K_{d})$ is a genuine fixed point. This presents a distinct scenario from real disorder cases, where the leading-order fixed point can be disrupted by higher-order fluctuations  Nattermann and Renz (1989). In future, it would be interesting to explore this new phase appearing in 3D.
@@
where $\hat{H}_{0}=\sum_{\mathbf{x},\mathbf{x}^{\prime}}h_{\mathbf{x},\mathbf{x}^{\prime}}\hat{a}^{\dagger}_{\mathbf{x}}\hat{a}_{\mathbf{x}^{\prime}}$ is a Hermitian Hamiltonian, with $\hat{a}^{\dagger}_{\mathbf{x}}(\hat{a}_{\mathbf{x}})$ being the creation (annihilation) operator at lattice site $\mathbf{x}$, and $\mathcal{D}[\hat{a}_{\mathbf{x}}]\hat{\rho}=\hat{a}_{\mathbf{x}}\hat{\rho}\hat{a}_{\mathbf{x}}^{\dagger}-\frac{1}{2}(\hat{a}_{\mathbf{x}}^{\dagger}\hat{a}_{\mathbf{x}}\hat{\rho}+\hat{\rho}\hat{a}_{\mathbf{x}}\hat{a}_{\mathbf{x}})$. On the right hand side of Eq. (6), the first term describes an unitary evolution given by $\hat{H}_{0}$, and the second term accounts local losses with site-dependent random rate $\gamma_{\mathbf{x}}$. Using Eq. (6), we can show that two-point correlation function $G_{\mathbf{x}_{1},\mathbf{x}_{2}}=\text{Tr}[\hat{\rho}\hat{a}_{\mathbf{x}_{1}}^{\dagger}\hat{a}_{\mathbf{x}_{2}}]$ satisfies the following matrix equation Sup
@@
where $G=\sum_{\mathbf{x},\mathbf{x}^{\prime}}G_{\mathbf{x},\mathbf{x}^{\prime}}|\mathbf{x}\rangle\langle\mathbf{x}^{\prime}|$, and $H=h^{T}+iV$ describes a Hermitian system $h^{T}$ subject to a random imaginary disorder $iV$ with $V_{\mathbf{x},\mathbf{x}^{\prime}}=\gamma_{\mathbf{x}}\delta_{\mathbf{x},\mathbf{x}^{\prime}}$. The equation is solved by $G(t)=e^{iHt}G(0)e^{-iH^{\dagger}t}$. Considering $N$ bosonic particles (e.g., photons in cavity) initially locate at $\mathbf{x}=\boldsymbol{0}$, i.e., $G_{\mathbf{x},\mathbf{x}^{\prime}}(0)=N\delta_{\mathbf{x},\mathbf{x}^{\prime}}\delta_{\mathbf{x},\boldsymbol{0}}$, the correlation matrix at time $t$ evolves to $G(t)=N|\phi(t)\rangle\langle\phi(t)|$ where $|\phi(t)\rangle=e^{iHt}|\boldsymbol{0}\rangle$ acts as a “quantum state” dictated by the “Hamiltonian” $H$. The probability distribution of surviving particles is given by $P(\mathbf{x},t)=G_{\mathbf{x},\mathbf{x}}(t)/\sum_{\mathbf{x}}G_{\mathbf{x},\mathbf{x}}(t)=|\langle\mathbf{x}|\phi(t)\rangle|^{2}/\langle\phi(t)|\phi(t)\rangle$. This coincides exactly with the particle evolution before, indicating the same dynamical scaling behavior in such an open system. Furthermore, similar dynamic behaviors can emerge from purely dissipative systems ($\hat{H}_{0}=0$) by considering suitable dissipators involving adjacent sites Sup ; Longhi (2023).
@@
Conclusion. In systems featuring complex disorders, the non-unitary time evolution facilitates a particle’s jumpy motion even amidst eigenstate localization, in sharp contrast to the conventional Anderson localization. Employing an intuitive optimization approach, we unveiled the universal dynamical scaling in this unconventional non-Hermitian transport, distinctly different from typical diffusive or ballistic transport. More interestingly, we find a close connection between the scaling behavior and the imaginary-part density of states (iDOS), especially the long-time behavior is dictated by the tail of iDOS. Our findings underscore the fundamental distinction between non-Hermitian and Hermitian localization, particularly in terms of dynamics. Numerous open questions persist, including the theoretical prediction and experimental observation of dynamic behaviors arising from different types of disorders, particularly in the regime with comparable kinetic and potential terms. It is also interesting to explore the combined dynamical consequence of Anderson localization and non-Hermitian skin effect Yao and Wang (2018); Kunst et al. (2018); Xiao et al. (2020); Helbig et al. (2020); Martinez Alvarez et al. (2018); Lee and Thomale (2019); Ashida et al. (2020); Bergholtz et al. (2021); Gohsrich et al. ; Jiang et al. (2019); Weidemann et al. (2022); Zeng et al. (2020); Wang et al. (2024). Additionally, the interplay between interactions and non-Hermitian disorders poses an intriguing avenue for exploration. On the experimental front, photonic systems provide a promising platform, as demonstrated by relevant experiments Weidemann et al. (2021). Furthermore, open systems featuring local random losses, such as dissipative cavity arrays, offer additional opportunities to validate our theory.
@@
Acknowledgements. This work is supported by NSFC under Grant No. 12125405 and National Key R&D Program of China (No. 2023YFA1406702). C.C. acknowledges support from NSFC under Grant No. 12347107.

### v2

@@ Abstract
where $g=v\eta^{2}/D^{4}$, $K_{d}=S_{d}/(2\pi)^{d}$ with $S_{d}$ being the solid angle in $d$ dimension, and the rescaling $\mathbf{x}\rightarrow e^{l}\mathbf{x}$ is performed. The flow equation yields two fixed points, $g_{1}^{\ast}=0$, $g^{\ast}_{2}=d(d-4)/[2(12-5d)K_{d}]$. Notably, $g_{1}^{\ast}$ is unstable for $d<4$, while $g^{\ast}_{2}$ is physically untenable for $d<12/5$ due to its negative value, conflicting with the definition. Thus, in 1D and 2D, the system does not transit to new phases upon rescaling, validating our analysis for weak disorders. However, for $d=3$, $g^{\ast}_{2}=1/(2K_{d})$ is a genuine fixed point, signifying a delocalization transition. In the delocalized phase with weak disorder, the system’s long-time behavior is governed by this fixed point, meaning the space-time relation remains invariant under rescaling. Consequently, the system exhibits dynamical scaling $|\mathbf{x}|\sim t^{3/5}$ Sup . This result is independent of the disorder statistics and beyond the approach outlined in Eqs. (2) and (3) which focuses on the eigenstate-localized regime.
@@
where $\hat{H}_{0}=\sum_{\mathbf{x},\mathbf{x}^{\prime}}h_{\mathbf{x},\mathbf{x}^{\prime}}\hat{a}^{\dagger}_{\mathbf{x}}\hat{a}_{\mathbf{x}^{\prime}}$ is a Hermitian Hamiltonian, with $\hat{a}^{\dagger}_{\mathbf{x}}(\hat{a}_{\mathbf{x}})$ being the creation (annihilation) operator at lattice site $\mathbf{x}$, and $\mathcal{D}[\hat{a}_{\mathbf{x}}]\hat{\rho}=\hat{a}_{\mathbf{x}}\hat{\rho}\hat{a}_{\mathbf{x}}^{\dagger}-\frac{1}{2}(\hat{a}_{\mathbf{x}}^{\dagger}\hat{a}_{\mathbf{x}}\hat{\rho}+\hat{\rho}\hat{a}^{\dagger}_{\mathbf{x}}\hat{a}_{\mathbf{x}})$. On the right hand side of Eq. (7), the first term describes a unitary evolution given by $\hat{H}_{0}$, and the second term accounts local losses with site-dependent random rate $\gamma_{\mathbf{x}}$. Using Eq. (7), we can show that two-point correlation function $G_{\mathbf{x}_{1},\mathbf{x}_{2}}=\text{Tr}[\hat{\rho}\hat{a}_{\mathbf{x}_{1}}^{\dagger}\hat{a}_{\mathbf{x}_{2}}]$ satisfies the following matrix equation Sup
@@
where $G=\sum_{\mathbf{x},\mathbf{x}^{\prime}}G_{\mathbf{x},\mathbf{x}^{\prime}}|\mathbf{x}\rangle\langle\mathbf{x}^{\prime}|$, and $H=h^{T}+iV$ describes a Hermitian system $h^{T}$ subject to a random imaginary disorder $iV$ with $V_{\mathbf{x},\mathbf{x}^{\prime}}=\gamma_{\mathbf{x}}\delta_{\mathbf{x},\mathbf{x}^{\prime}}$. The equation is solved by $G(t)=e^{iHt}G(0)e^{-iH^{\dagger}t}$. Considering $N$ bosonic particles (e.g., photons in a cavity) initially locate at $\mathbf{x}=\boldsymbol{0}$, i.e., $G_{\mathbf{x},\mathbf{x}^{\prime}}(0)=N\delta_{\mathbf{x},\mathbf{x}^{\prime}}\delta_{\mathbf{x},\boldsymbol{0}}$, the correlation matrix at time $t$ evolves to $G(t)=N|\phi(t)\rangle\langle\phi(t)|$ where $|\phi(t)\rangle=e^{iHt}|\boldsymbol{0}\rangle$ acts as a “quantum state” dictated by the “Hamiltonian” $H$. The probability distribution of surviving particles is given by $P(\mathbf{x},t)=G_{\mathbf{x},\mathbf{x}}(t)/\sum_{\mathbf{x}}G_{\mathbf{x},\mathbf{x}}(t)=|\langle\mathbf{x}|\phi(t)\rangle|^{2}/\langle\phi(t)|\phi(t)\rangle$. This coincides exactly with the particle evolution before, indicating the same dynamical scaling behavior in such an open system. Furthermore, similar dynamic behaviors can emerge from purely dissipative systems ($\hat{H}_{0}=0$) by considering suitable dissipators involving adjacent sites Sup ; Longhi (2023).
@@
Conclusion. In systems featuring complex disorders, the non-unitary time evolution facilitates a particle’s jumpy motion even amidst eigenstate localization, in sharp contrast to the conventional Anderson localization. Employing an intuitive optimization approach, we unveiled the universal dynamical scaling in this unconventional non-Hermitian transport, distinctly different from typical diffusive or ballistic transport. More interestingly, we find a close connection between the scaling behavior and the imaginary-part density of states (ImDOS), especially the long-time behavior is dictated by the tail of ImDOS. Our findings underscore the fundamental distinction between non-Hermitian and Hermitian localization, particularly in terms of dynamics. Numerous open questions persist, including the theoretical prediction and experimental observation of dynamic behaviors arising from different types of disorders, particularly in the regime with comparable kinetic and potential terms. It is also interesting to explore the combined dynamical consequence of Anderson localization and non-Hermitian skin effect Yao and Wang (2018); Yao et al. (2018); Kunst et al. (2018); Xiao et al. (2020); Helbig et al. (2020); Martinez Alvarez et al. (2018); Lee and Thomale (2019); Ashida et al. (2020); Bergholtz et al. (2021); Gohsrich et al. ; Jiang et al. (2019); Weidemann et al. (2022); Zeng et al. (2020); Wang et al. (2025b). Additionally, the interplay between interactions and non-Hermitian disorders poses an intriguing avenue for exploration. On the experimental front, photonic systems provide a promising platform, as demonstrated by relevant experiments Weidemann et al. (2021). Furthermore, open systems featuring local random losses, such as dissipative cavity arrays, offer additional opportunities to validate our theory.
@@
Acknowledgements. This work is supported by National Key R&D Program of China (No. 2023YFA1406702) and NSFC under Grant No. 12125405. B. L. is partially supported by NSFC under Grant No. 12404185. C.C. acknowledges support from NSFC under Grant No. 12347107.